Shift report 20011009 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Oct 09 2001 - 08:22:27 EDT

  • Next message: Flemming Videbaek: "Fw: Security at BNL"

    ------------------------------------------------------------
    Date: Tue Oct  9 08:22:27 2001, Shift: 00:00-08:00
    Supervisor: Michael Murray, with: none
    Log book#9 pages 123 - 128
    
    Shift Summary:
    ------------------------------------------------------------
    1) One long store from 00:30 until 4:30. PHENIX and BRAHMS
    had voted to end the store at 3.30 but there was no 
    gold in the tandum. At the begging of the store RHIC steered
    in other IRs. I asked them to steer in BRAHMS but got little
    improvement. As of 8.10am beam is stored and cogged
    and RHIC promises to steer in all IRs.
    2) Took one useful run 5259 with 300K triggers and one
    not so useful one 5238 without the TPCs. (my apologies)
    Since the current in T3,T4 and T5 was low FV suggested
    raising their anode voltages by 20V. They did not trip 
    until 4.30 when RHIC suddendenly dumped the beam.
    3) No detector problems but tape drive in small room was 
    beeping most of the night. Betty will have a look at it.
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Oct 09 2001 - 08:24:44 EDT